# Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation

### v1

@@ Abstract
Large language models (LLMs) with instruction following capabilities have demonstrated impressive problem-solving abilities. While synthesizing instructional data from unsupervised text has become a common approach for training such models, conventional methods rely heavily on human effort for data annotation. Although existing automated synthesis paradigms have alleviated this constraint, they still exhibit significant limitations in ensuring adequate diversity and difficulty of synthesized instructions. To address these challenges, we propose Self-Foveate, an innovative LLM-driven method for instruction synthesis. This approach introduces a "Micro-Scatter-Macro" multi-level foveation methodology that effectively guides the LLM to deeply excavate fine-grained information embedded in unsupervised text, thereby enhancing both the diversity and difficulty of synthesized instructions. Comprehensive experiments across multiple unsupervised corpora and diverse model architectures validate the effectiveness and superiority of our proposed method. We publicly release our data and codes: https://github.com/Mubuky/Self-Foveate
@@
Figure 1: Illustration of (a) Self-Foveate in contrast with (b) Baseline Self-QA. For Self-Foveate, the Multi-Level Foveation enables the LLM to extract details (highlighted in distinct colors) of the text, subsequently synthesizing instructions with diversity and difficulty via distinct synthesis paradigms. In comparison, Self-QA employs single-step generation that produces instruction candidates with simplicity and monotony.
Related Data w/o Human Correctness Diversity Difficulty
Work Utilization Annotation Guarantee Augment Augment
Self-Instruct (Wang et al., 2023a) Seed QA examples $\checkmark$
Self-Align (Sun et al., 2023) Seed QA examples $\checkmark$ $\checkmark$
Self-Chat (Xu et al., 2023) Dialogue $\checkmark$
Self-QA (Zhang and Yang, 2023) Unsupervised Knowledge $\checkmark$ $\checkmark$ $\checkmark$
LongForm (Köksal et al., 2024) Web dataset $\checkmark$ $\checkmark$ $\checkmark$
Humpback (Li et al., 2024b) Web dataset $\checkmark$ $\checkmark$
ISARA (Guo et al., 2024) Seed QA examples $\checkmark$ $\checkmark$ $\checkmark$
Wiki2023 (Jiang et al., 2024) Unsupervised Text $\checkmark$ $\checkmark$
Self-Foveate (Ours) Unsupervised Text $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 1 Introduction
Large language models (LLMs), such as GPT-4o (OpenAI, 2024), Claude 3.5 Sonnet (Anthropic, 2024), and Llama 3.1 (Meta, 2024), have garnered significant attention due to their exceptional instruction-following capabilities (Zhou et al., 2023b), with their continuously enhanced problem-solving abilities (Cobbe et al., 2021; Dua et al., 2019) being increasingly recognized. A critical component in training such models typically involves fine-tuning with extensive supervised question-answering instruction data (Wang et al., 2023b). However, substantial challenges persist in constructing large-scale, high-quality supervised fine-tuning (SFT) data for instruction tuning (Zhang et al., 2024). The prohibitive costs associated with human annotation (Wang et al., 2023a; Sun et al., 2023; Xu et al., 2023), coupled with difficulties in ensuring data diversity and quality control (Ge et al., 2024), continue to impede technological advancements. Given the demonstrated excellence and ongoing improvements in LLMs’ instruction-following and generative capabilities, researchers are actively exploring effective methodologies to leverage these models for synthetic data generation (Wang et al., 2023a; Zhang and Yang, 2023; Nayak et al., 2024b; Wu et al., 2024). The primary objectives are to produce high-quality, cost-efficient datasets that reduce reliance on expensive human annotation (Ling et al., 2024), while enhancing the diversity and difficulty of automatically synthesized instructions to improve the performance of fine-tuned models on downstream tasks.
@@
A promising recent paradigm in synthetic instruction synthesis is unsupervised text-based instruction synthesis. The advantages of this paradigm become particularly pronounced when handling massive unsupervised text corpora – a ubiquitous resource containing rich world knowledge and linguistic patterns. By leveraging LLMs’ intrinsic capabilities in contextual understanding and logical reasoning, this paradigm eliminates the need for manual annotation while confining the scope of instruction synthesis to the given unsupervised textual materials. The pioneering work by Zhang and Yang (2023) has demonstrated the feasibility of extracting instructions from unsupervised textual data.
@@
Despite advancements in synthesizing instruction data from unsupervised text and the proposal of automated LLM-based methods like Self-QA (Zhang and Yang, 2023) to mitigate these issues – approaches that have significantly optimized the data generation pipeline while reducing human labor costs – as shown in Table 1, the following challenges remain unresolved: (1) Diversity of Instructions: While existing frameworks continuously refine data generation strategies and enhance post-synthesis filtering for instruction data, limitations persist in the diversity of synthesized instructions. Models trained on such synthetic datasets frequently exhibit insufficient generalization capabilities and may even suffer from performance degradation. (2) Difficulty of Instructions: Current synthesis methodologies generally lack emphasis on instruction complexity and depth. For instance, Self-QA (Zhang and Yang, 2023) directly acquires instructions through single-step generation without guaranteed difficulty levels (as illustrated in Figure 1). Synthesized instructions often demonstrate simplistic structures and inadequate comprehension of inter-entity relationships, failing to effectively stimulate models to produce nuanced responses. This deficiency becomes particularly pronounced when handling complex queries or generating contextually rich responses. The absence of high-difficulty instruction design constrains model performance in real-world applications requiring advanced cognitive engagement and problem-solving capabilities. This collective evidence indicates that unsupervised text data still harbors substantial untapped potential for instruction synthesis.
@@
To address these challenges, we attempt to leverage unsupervised text itself, observing that textual data inherently contains abundant detailed information encompassing entities (e.g., QuantumX processor), attributes, relations, and writing techniques – such as the metaphorical comparison of “speed” to “not just a luxury” in the statement “In the realm of technology, speed is not just a luxury”, which implicitly analogizes speed as a fundamental necessity through subtle analogy, thereby emphasizing its critical importance (as multi-color annotated in Figure 1). This information remains underutilized in existing methods like Self-QA.
@@
Motivated by these observations, this paper proposes Self-Foveate, a comprehensive LLM-based methodology designed to automatically synthesize instructions from unsupervised text. Diverging from prior research, as illustrated in Figure 1, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology to comprehensively excavate detailed information from raw text, subsequently synthesizing instructions with enhanced diversity and difficulty through three synthesis paradigms. Furthermore, Self-Foveate integrates a data regeneration module to improve the fidelity and quality of instructions to source text.
@@
    We focus on unsupervised text-based instruction synthesis tasks, revealing limitations in existing works regarding diversity and difficulty.
@@
    We propose Self-Foveate, a method that synthesizes instructions with diversity and difficulty through LLMs based on unsupervised text.
@@
    We conducted extensive experiments to evaluate Self-Foveate, covering diversity and difficulty analysis, downstream task capabilities, and data scale trend analysis. The results demonstrate Self-Foveate’s superiority in unsupervised text-based instruction synthesis and suggest promising directions for future research.
@@ 2 Related Work
Instruction Tuning
Multitask instruction fine-tuning (Wei et al., 2022) of language models significantly enhances their ability to follow instructions and generalize to new unseen tasks (Sanh et al., 2022; Mishra et al., 2022; Chung et al., 2022; Longpre et al., 2023; Zhou et al., 2023a; Li et al., 2024b). In our work, we utilize data from unsupervised text synthesis to conduct instruction tuning, enabling the model to better adapt to specific domain tasks.
Synthetic Data Generation
LLMs have showcased remarkable capabilities in data synthesis (Long et al., 2024), facilitating the creation of extensive synthetic datasets for pretraining and fine-tuning, thereby progressively supplanting labor-intensive manual data scraping and selection (Liu et al., 2024), and mitigating the constraints that data imposes on model capability growth (Villalobos et al., 2024).
Distinct from earlier approaches centered on traditional language models (Schick and Schütze, 2021), LLMs present enhanced potential for generating high-quality synthetic data across diverse applications, including online translation (Oh et al., 2023), named entity recognition (Xiao et al., 2023), benchmark creation (Wang et al., 2024; Wei et al., 2024), and data diversity enhancement (Dai et al., 2023; Chung et al., 2023; Hong et al., 2024).
The concept of synthetic input-output pairs for instruction tuning advances by requiring that the data generated by LLMs be diverse, accurate, and difficult, often leveraging LLMs on a set of seed task demonstrations or user-provided unsupervised text to create new synthetic tasks (Wang et al., 2023a; Honovich et al., 2023; Zhang and Yang, 2023; Taori et al., 2023; Peng et al., 2023; Yuan et al., 2024; Li et al., 2024a).
Our work advances the generation of synthetic input-output pairs by developing a paradigm that integrates multi-level foveation into the creation of instruction tuning datasets, forming a comprehensive framework without any human annotation.
@@ 3 Self-Foveate
In this section, we will introduce the proposed Self-Foveate, denoted as $\mathcal{F}$, a multi-level automated method for synthesizing instructions that leverages only unsupervised text without the need for human-annotated samples. Self-Foveate consists of three levels and a re-synthesis module. Formally, consider an original unsupervised text set $\mathcal{D}$, the proposed method Self-Foveate operates by transforming each element $d_{i}$ from $\mathcal{D}$ through a multi-level synthesis path $\mathcal{F}_{j}$. In practice, Self-Foveate is an operation applied independently to each document $d_{i}$ ($d_{i}\in\mathcal{D}$). The final generated dataset, $\mathcal{D}_{\text{gen}}$, is accumulated from the data subsets generated by all elements:
$\mathcal{D}_{\text{gen}}=\mathcal{F}(\mathcal{D})=\sum_{d_{i}\in\mathcal{D}}\sum_{\mathcal{F}_{j}\in\mathcal{F}}\mathcal{F}_{j}(d_{i})$ (1)
During the synthesis process, the objectives of Self-Foveate emphasize maximizing the diversity and difficulty of the generated instructions.
To ensure optimal functionality, Self-Foveate incorporates three levels and one module as shown in Figure 2:
@@
    Micro-foveate Level: This level captures the essential content within unsupervised text and acquires instructions through reverse synthesis. It plays a crucial role in maintaining the focus of synthesized instructions on the significant information in the text.
@@
    Scatter-foveate Level: This level aims to combine key information scattered throughout the unsupervised text and uses direct synthesis to generate instructions. It stimulates LLMs to synthesize instructions that possess a profound understanding of the relationships between different contents in the text.
@@
    Macro-foveate Level: This level focuses on and transcribes larger-grained information within the text. It plays a key role in enhancing the instructions’ deep comprehension of the overall information in the text.
@@
    Re-synthesis Module: This module is used to perform post-synthesis filtering on the generated instructions, excluding outliers in the instruction synthesis process and conducting reference-based re-synthesis for abnormally generated instructions to reduce information gaps in the instruction set and improve synthesis success rates.
@@ 3.1 Micro-foveate Level
General unsupervised text typically contains several primary entity types, a greater number of secondary entity types, and attributes of all these entity types. However, instruction sets directly synthesized by teacher LLMs may struggle to intentionally and comprehensively cover all the important content within such unsupervised text. This could lead to the synthesized instruction sets missing critical information from the original text, thereby resulting in weaker downstream model performance. To address this, we employ the micro-foveate mechanism and reverse synthesis to ensure that these key pieces of information are preserved in the synthesized instruction sets. Attention to a broader range of information enhances the diversity of the synthesized instruction sets, while the exploration of entity attributes also increases the difficulty.
@@
This mechanism aims to guide the teacher LLM to synthesize instructions from a fine-grained perspective of the unsupervised text. We introduce the concept of “foveate elements”, which broadly encompass all entities and their attributes within a given text. We guide the large model to extract more foveate elements than needed and then select a certain number of foveate elements with the highest cosine similarity between the text and the embedding of each foveate element.
@@
Based on the selected foveate elements, we employ the reverse synthesis method to generate instructions. Specifically, each foveate element is treated as a potential answer to an instruction, guiding the teacher LLM to synthesize instructions from the unsupervised text and then re-synthesize the answers. Even though the foveate elements can already serve as answers or parts of answers, to enhance the fluency, completeness, and accuracy of the answers, we choose to regenerate the answers in the reverse synthesis step rather than directly using the foveate elements as answers.

### v2

@@ Abstract
Synthesizing high-quality instruction data from unsupervised text is a promising paradigm for training large language models (LLMs), yet automated methods for this task still exhibit significant limitations in the diversity and difficulty of synthesized instructions. To address these challenges, we propose Self-Foveate, an LLM-driven method for instruction synthesis. Inspired by hierarchical human visual perception, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology that guides the extraction of textual information at three complementary granularities, from fine-grained details through cross-region connections to holistic patterns, thereby enhancing both the diversity and difficulty of synthesized instructions. Furthermore, a re-synthesis module is incorporated to improve the fidelity of instructions to source text and their overall quality. Comprehensive experiments across multiple unsupervised corpora and diverse model architectures demonstrate that Self-Foveate consistently outperforms existing methods. We publicly release our code at https://github.com/Mubuky/Self-Foveate
@@
Figure 1: Illustration of (a) Self-Foveate in contrast with (b) Baseline Self-QA. For Self-Foveate, the multi-level foveation enables the LLM to extract details (highlighted in distinct colors) of the text, subsequently synthesizing instructions with diversity and difficulty via distinct synthesis paradigms. In comparison, Self-QA employs single-step generation that produces simple and monotonous instruction candidates.
Related Data w/o Human Correctness Diversity Difficulty Work Utilization Annotation Guarantee Augment Augment Self-Instruct (Wang et al., 2023a) Seed QA examples $\checkmark$ Self-Align (Sun et al., 2023) Seed QA examples $\checkmark$ $\checkmark$ Self-Chat (Xu et al., 2023) Dialogue $\checkmark$ Self-QA (Zhang and Yang, 2023) Unsupervised Knowledge $\checkmark$ $\checkmark$ $\checkmark$ LongForm (Köksal et al., 2024) Web dataset $\checkmark$ $\checkmark$ $\checkmark$ Humpback (Li et al., 2024b) Web dataset $\checkmark$ $\checkmark$ ISARA (Guo et al., 2024) Seed QA examples $\checkmark$ $\checkmark$ $\checkmark$ Wiki2023 (Jiang et al., 2024) Unsupervised Text $\checkmark$ $\checkmark$ Self-Foveate (Ours) Unsupervised Text $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 1 Introduction
Large language models (LLMs), such as GPT-4o (OpenAI et al., 2024), Claude 3.5 Sonnet (Anthropic, 2024), and Llama 3.1 (Grattafiori et al., 2024), have demonstrated exceptional instruction-following capabilities (Zhou et al., 2023b) and increasingly strong problem-solving abilities (Cobbe et al., 2021; Dua et al., 2019). A critical component in training such models involves fine-tuning with extensive supervised question-answering instruction data (Wang et al., 2023b). However, substantial challenges persist in constructing large-scale, high-quality supervised fine-tuning (SFT) data for instruction tuning (Zhang et al., 2024). The prohibitive costs of human annotation (Wang et al., 2023a; Sun et al., 2023; Xu et al., 2023), coupled with difficulties in ensuring data diversity and quality control (Ge et al., 2024), continue to impede technological advancements.
Given the strong instruction-following and generative capabilities of modern LLMs, researchers are actively exploring methodologies to leverage these models for synthetic data synthesis (Wang et al., 2023a; Zhang and Yang, 2023; Nayak et al., 2024b; Wu et al., 2024). The primary objectives are to produce high-quality datasets with enhanced diversity and difficulty, thereby improving the performance of fine-tuned models on downstream tasks.
@@
A promising recent paradigm in instruction data synthesis is unsupervised text-based instruction synthesis. The advantages of this paradigm become particularly pronounced when handling massive unsupervised text corpora, a ubiquitous resource containing rich world knowledge and linguistic patterns. By leveraging LLMs’ intrinsic capabilities in contextual understanding and logical reasoning, this paradigm eliminates the need for manual annotation while grounding instruction synthesis in the given unsupervised textual materials.
@@
Recent automated methods such as Self-QA (Zhang and Yang, 2023) have significantly optimized the data generation pipeline while reducing human labor costs. However, as shown in Table 1, critical challenges remain unresolved: (1) Diversity of Instructions: While existing frameworks continuously refine data generation strategies and enhance post-synthesis filtering for instruction data, limitations persist in the diversity of synthesized instructions, which tend to exhibit repetitive patterns in both structure and topic coverage, potentially constraining the generalization of fine-tuned models. (2) Difficulty of Instructions: Current synthesis methodologies generally lack emphasis on instruction complexity and depth. For instance, Self-QA (Zhang and Yang, 2023) directly acquires instructions through single-step generation without guaranteed difficulty levels (as illustrated in Figure 1). Synthesized instructions often exhibit simplistic structures and fail to capture complex inter-entity relationships, limiting models’ ability to produce nuanced responses. These limitations suggest that existing methods fail to fully exploit the rich, multi-granularity information inherent in unsupervised text, leaving substantial potential untapped.
@@
To address these challenges, we observe that textual data inherently contains abundant detailed information encompassing entities (e.g., QuantumX processor), attributes, relations, and writing techniques. For instance, the statement “In the realm of technology, speed is not just a luxury” implicitly frames speed as a fundamental necessity through subtle analogy (as multi-color annotated in Figure 1). This rich, multi-granularity information remains underutilized in existing methods.
This multi-level nature of textual information finds a compelling parallel in human visual perception. In the human visual system, the fovea centralis, a small region of the retina with the highest density of photoreceptor cells, enables sharp central vision for resolving fine details at the point of fixation (Wandell, 1995). Beyond this focal point, rapid saccadic eye movements shift the fovea across different regions of interest, and the brain integrates information from these dispersed fixation points to comprehend spatial relationships (Findlay and Gilchrist, 2003). Meanwhile, peripheral vision captures the global structure and holistic patterns of the scene at a coarser resolution. This multi-level visual processing, spanning fine-grained foveal detail, cross-region saccadic integration, and holistic peripheral awareness, inspires our approach to multi-granularity text comprehension.
@@
Motivated by these textual observations and the analogous multi-level processing in human foveation, this paper proposes Self-Foveate, a comprehensive LLM-based methodology designed to automatically synthesize instructions from unsupervised text. As illustrated in Figure 1, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology in which micro-foveate mirrors foveal fixation on fine-grained details, scatter-foveate emulates saccadic integration across dispersed information, and macro-foveate parallels peripheral perception of holistic patterns, thereby systematically extracting detailed information from raw text and subsequently synthesizing instructions with enhanced diversity and difficulty through three synthesis paradigms. Furthermore, Self-Foveate integrates a re-synthesis module to improve both the fidelity of instructions to source text and their overall quality.
@@
    We systematically analyze unsupervised text-based instruction synthesis, identifying critical limitations of existing methods in both instruction diversity and difficulty.
@@
    We propose Self-Foveate, a multi-level foveation methodology inspired by human visual perception that synthesizes diverse and challenging instructions from unsupervised text via three complementary strategies: micro-, scatter-, and macro-foveation.
@@
    We conduct extensive experiments covering diversity and difficulty analysis, downstream task evaluation, and data scaling analysis. Results demonstrate that Self-Foveate consistently outperforms existing methods in unsupervised text-based instruction synthesis.
@@ 2 Related Work
Instruction Data Synthesis
Multitask instruction fine-tuning (Wei et al., 2022) substantially improves language models’ ability to follow instructions and generalize to unseen tasks (Sanh et al., 2022; Mishra et al., 2022; Chung et al., 2022; Longpre et al., 2023; Zhou et al., 2023a; Li et al., 2024b). As manually curating large-scale instruction data remains costly (Liu et al., 2024) and the constraints that data availability imposes on model capability continue to grow (Villalobos et al., 2024; Long et al., 2024), recent work leverages LLMs to synthesize instruction–response pairs automatically. Self-Instruct (Wang et al., 2023a) bootstraps new tasks from a small set of seed demonstrations; Unnatural Instructions (Honovich et al., 2023) and Alpaca (Taori et al., 2023) scale this idea with minimal human intervention; and Self-QA (Zhang and Yang, 2023) further removes the need for seed tasks by grounding synthesis in unsupervised text. Other efforts refine this pipeline through GPT-4 distillation (Peng et al., 2023), self-rewarding mechanisms (Yuan et al., 2024), and generalized synthesis frameworks (Li et al., 2024a). Our work follows the unsupervised text-based synthesis paradigm, focusing specifically on the quality dimensions of diversity and difficulty that prior pipelines leave underaddressed.
Diversity and Difficulty of Synthesized Instructions
Despite the rapid progress in instruction data synthesis, two persistent challenges remain underexplored. First, synthesized instructions often lack variety in both form and subject matter, narrowing the diversity of the resulting datasets (Chung et al., 2023; Ge et al., 2024). Second, most synthesis pipelines lack explicit mechanisms to control instruction difficulty; for example, Self-QA (Zhang and Yang, 2023) produces instructions in a single pass, offering no explicit control over their complexity. While earlier approaches based on smaller pretrained language models (Schick and Schütze, 2021) were constrained by limited model capacity, even modern LLM-based pipelines have not fully addressed these quality dimensions. A common limitation is that existing methods treat source text as a monolithic input rather than systematically exploiting its multi-granularity structure, leaving substantial room for more varied and challenging instructions. To close this gap, Self-Foveate introduces a multi-level foveation methodology that guides the extraction of textual information at three complementary granularities and yields instructions of greater variety and complexity, all without human annotation.
@@ 3 Self-Foveate
We propose Self-Foveate, denoted as $\mathcal{F}$, a multi-level method for synthesizing instructions from unsupervised text without human-annotated samples. Self-Foveate consists of three levels and a re-synthesis module. Formally, given an unsupervised text set $\mathcal{D}$, Self-Foveate transforms each document $d_{i}\in\mathcal{D}$ through a set of synthesis paths $\{\mathcal{F}_{j}\}$. The final synthesized dataset, $\mathcal{D}_{\text{gen}}$, is the union of subsets produced by applying every synthesis path to every document:
$\mathcal{D}_{\text{gen}}=\mathcal{F}(\mathcal{D})=\bigcup_{d_{i}\in\mathcal{D}}\bigcup_{\mathcal{F}_{j}\in\mathcal{F}}\mathcal{F}_{j}(d_{i})$ (1)
The synthesis process is designed to maximize both the diversity and difficulty of the resulting instructions.
As shown in Figure 2, Self-Foveate incorporates three levels and one module:
@@
    Micro-foveate Level: Extracts fine-grained foveate elements (entities and attributes) from unsupervised text, then uses each element as a seed answer to synthesize instructions via reverse synthesis (§3.1).
@@
    Scatter-foveate Level: Broadly extracts foveate elements and combines them into foveate groups, then applies direct synthesis to produce instructions requiring cross-entity reasoning (§3.2).
@@
    Macro-foveate Level: Identifies text segments employing rhetorical or figurative devices as foveate segments, then converts them into instructional form via transcription synthesis (§3.3).
@@
    Re-synthesis Module: Filters the synthesized instruction set for unanswerable instructions and iteratively re-synthesizes them using successfully synthesized examples as references, improving fidelity to the source text (§3.4).
@@ 3.1 Micro-foveate Level
Unsupervised text typically contains a hierarchy of information: primary entities, secondary entities, and their respective attributes. Without explicit guidance, a teacher LLM synthesizing instructions from such text tends to focus on salient surface-level content, overlooking secondary entities and fine-grained attributes. The resulting instruction sets therefore miss critical information, limiting downstream model performance. The micro-foveate level addresses this gap through a fine-grained extraction mechanism paired with reverse synthesis, ensuring that individually important pieces of information are preserved in the synthesized instruction sets. Covering a broader range of entities and attributes enhances diversity, while probing specific entity attributes increases difficulty.
@@
This mechanism aims to guide the teacher LLM to synthesize instructions from a fine-grained perspective of the unsupervised text. We introduce the concept of “foveate elements”, which broadly encompass all entities and their attributes within a given text. We guide the teacher LLM to extract more foveate elements than needed and then retain those whose embeddings have the highest cosine similarity to the full text embedding, filtering out extractions that are tangential to the source content.
@@
Based on the selected foveate elements, we employ the reverse synthesis method to produce instructions. Specifically, we treat each foveate element as a potential answer to an instruction and guide the teacher LLM to synthesize a corresponding instruction from the unsupervised text, then derive the final answer. Even though the foveate elements can already serve as answers or parts of answers, to enhance the fluency, completeness, and accuracy of the answers, we choose to derive the answers anew in the reverse synthesis step rather than directly using the foveate elements as answers.